# Joint Pedestrian Trajectory Prediction through Posterior Sampling

### v1

@@ III-E Trajectory Latent Representation
Inspired by the advances in image synthesis bought by latent diffusion models [25] and the successful application of latent representation in trajectory prediction [26], we further consider transforming the spatial coordinates to latent space through Principal Component Analysis (PCA). PCA enables a low-rank approximation of trajectory data $X^{i}\in\mathbb{R}^{T*2}$ by mapping it to the latent representation $x^{i}\in\mathbb{R}^{k}$, which stands for linear combination of first-$k$ principal components. This transformation serves to mitigate the effects of noisy data, leading to a more refined and smoother trajectory. Meanwhile, the latent representation of trajectory contains its geometry and temporal characteristics, saving the need for additional temporal encoding steps in the network.
@@ IV Framework Architecture
As we re-formulate the task of prediction as trajectory inpainting, our proposed framework requires a pre-trained DDPM that models the joint distribution of full-length trajectories $p(x)$, which is not conditioned on history motion $c$. Through iteratively adding posterior guidance $-\lambda\nabla_{x_{t}}\mathcal{L}$ during inference time and extract the future part $y$ from generated $x$, we achieve the goal to sample from the conditional distribution $p(y|c)$. We argue that our framework is model-agnostic and that various reasonable network designs of the denoising module are feasible for our methodology. In this section, we present a specific implementation of a lightweight denoising module. We will discuss our data representation method, module architecture, and guidance design in the following.
Data Representation.   To achieve a rotation-invariant representation of agent motion, for each agent in the scene, we rotated and normalized their trajectory to their current position, decoupling the temporal dependency of history motion and social interaction. Parallelly, we extract the spatial relativity of agent trajectories by establishing a graph representation, where node $x^{i},i=1,\dots,N_{a}$ stands for agent motion and edge $E^{ij},i,j=1,\dots,N_{a}$ is characterized by a 6D vector describing their relative position at the current timestep: relative distance $d^{ij}$, direction vector $r^{ij}$ under the reference frame of agent $i$, and relative heading vector $h^{ij}=\{\theta^{ij},cos\theta^{ij},sin\theta^{ij}\}$.
@@
$\mathcal{L}_{rep}=weight\cdot\frac{1}{N_{a}}\sum_{i,j,k}\mathrm{max}\{(1-\frac{1}{r}dist^{ij}_{k}),0\},$ (15)
where $r$ is the repeller threshold and $dist^{ij}_{k}\in\mathbb{R}^{N_{a}\times N_{a}\times(T_{hist}+T_{fut})}$ is the distance between agent $i$ and $j$ at timestep $t$. Practically, the repeller loss serves as valuable prior knowledge in scenario generation.
@@ V Experiments
In this section, we exhibit the capacity and flexibility of our framework by adapting our model to four distinct tasks without retraining: Trajectory prediction, noisy data prediction, missing data prediction, and controllable generation. We present two models within our proposed framework: the foundational GFTD and its variant, GFTD-RePaint, which is specialized for the basic trajectory prediction task.
@@ V-B Trajectory Prediction and Controllable Generation
TABLE IV: Prediction with Missing Data
@@
25% Joint AgentFormer 2.000 / 3.352 0.610 / 1.043 1.674 / 2.805 1.832 / 3.279 2.093 / 3.092 1.642 / 2.714
Ours 0.528 / 0.912 0.181 / 0.311 0.662 / 1.326 0.342 / 0.677 0.314 / 0.632 0.405 / 0.772
50% Joint AgentFormer 1.763 / 3.007 0.661 / 1.090 1.746 / 2.953 1.887 / 3.379 1.989 / 2.928 1.609 / 2.671
Ours 0.514 / 0.843 0.186 / 0.316 0.703 / 1.390 0.343 / 0.673 0.323 / 0.643 0.414 / 0.773
75% Joint AgentFormer 1.973 / 3.200 0.688 / 1.118 1.936 / 3.347 1.966 / 3.500 1.731 / 2.567 1.659 / 2.746
Ours 0.493 / 0.775 0.241 / 0.423 0.859 / 1.648 0.377 / 0.726 0.419 /0.812 0.478 / 0.877
@@ V-C Prediction with Noisy History
To assess the robustness of our proposed method against noisy input, we perturbed the observed history trajectories by introducing random Gaussian noise $n\sim\mathcal{N}(0,\sigma I)$ of varying standard deviation $\sigma$. Specifically, we introduced two levels of perturbation: a slight level, characterized by a standard Gaussian noise with a 0.05-meter $\sigma$, and a heavy level, with a standard Gaussian noise with a 0.15-meter $\sigma$. Subsequently, we compared the performance of our model with the Joint AgentFormer baseline under identical experimental conditions.
As shown in TABLE LABEL:table:_noisy_data, we observed that our model performs better at heavy noise levels, providing trustworthy prediction even if the data is highly corrupted. The average JADE/JFDE performance of our method on 5 benchmarks is close to the baseline but only degenerated by $46.58\%/38.38\%$ at the heavy level. By contrast, the performance of the baseline method had degenerated by $57.08\%/50.36\%$ at heavy noise level. To reasonably state that the guided posterior sampling accounted for the robustness against noisy data, we further compared it with GFTD-RePaint. GFTD-RePaint includes gradually replacing the history part of the intermediate noisy sample by ground truth history observation, thus ensuring the generated samples were strictly conditioned on the noisy history data. The experimental results in Fig. 4 show that despite having a near performance to GFTD model under clean data conditions, the GFTD-RePaint performs much worse on noisy data, especially under heavy-level noise. This is in accordance with our anticipation, proving the superiority of our proposed framework at noisy conditions rooted in the design of ”soft conditioning” enabled by posterior sampling.
@@ V-D Prediction with Incomplete History
Through the incomplete data prediction task we show the adaptation capacity of our GFTD framework. By masking the initial $25\%$, $50\%$, and $75\%$ steps of history data, we mimic a practical situation where sensors are partially blocked by obstacles and only limited observation is available. We directly evaluated GFTD on these compromised data without retraining. Inferencing missing data could be done easily with no modification to the sampling process since the guidance only enforces conditions using available history steps. As displayed in TABLE LABEL:table:_missing_data, GFTD exhibits competitive performance under incomplete data settings. Even when $75\%$ of history observation is missing, the average JADE/JFDE performance only degenerated for $21.01\%/16.62\%$ and outperformed S-GAN with general data input by a large margin. In comparison, Joint AgentFormer completely failed to present plausible predictions when directly applied to missing data.
@@ VI Conclusion & Discussion
In this work, we present the Guided Full Trajectory Diffuser, a novel diffusion model-based framework for representing the joint distribution of full trajectories, converting trajectory prediction and controllable generation into a unified inverse problem. We formulate the prediction task as spatial-temporal inpainting, a general noisy inverse problem that can be solved through diffusion posterior sampling. Under this framework, the generated trajectories are not rigidly constrained by their historical observations; instead, we approximate such conditions by gradually adding posterior guidance. Such a unique design enables flexibility and adaptivity, resulting in not only competitive performance in joint trajectory prediction tasks but also in controllable generation, particularly in scenarios with noise injection or incomplete historical data. Moreover, our framework is compatible with plug-and-play modules including RePaint, thus expandable to task-oriented enhancements.

### v2

@@ III-E Trajectory Latent Representation
Inspired by the advances in image synthesis bought by latent diffusion models [28] and the successful application of low-rank latent representation in trajectory prediction [29] [23], we further consider transforming the spatial-temporal features of trajectory $X^{i}\in\mathbb{R}^{T*d}$ into latent features $x^{i}\in\mathbb{R}^{k}$, $k\ll T*d$ through Principal Component Analysis (PCA). This notation stands for the linear combination of the first principal components. This simple but efficient transformation serves to mitigate the effects of noisy data, leading to more consistent and smoother predicted trajectories. Meanwhile, the latent representation contains its geometry and temporal characteristics, saving the need for additional temporal encoder blocks.
@@ IV Framework Architecture
As we re-formulate the prediction task as trajectory inpainting, our proposed framework requires a pre-trained DDPM that models the joint distribution of full-length trajectories that is not conditioned on history motion $c$. Through iteratively adding posterior guidance $-\lambda\nabla_{x_{t}}\mathcal{L}$ during inference time and extracting the future parts $y$ from generated trajectories $x$, we achieve the goal to sample future trajectories conditioned on given histories. We argue that our framework is model-agnostic that various reasonable network designs of the denoising module are feasible for our methodology. This section presents a specific implementation of a lightweight denoising module. We will discuss our data representation method, module architecture, and guidance design in the following.
Data Representation.   To achieve a rotation-invariant representation of agent motion, for each agent in the scene, we normalized their trajectory according to their current positions and headings. Parallelly, we extract the spatial relativity of agents under current frame by establishing a graph representation, where node $x^{i},i=1,\dots,N_{a}$ stands for agent motion and edge $E^{ij},i,j=1,\dots,N_{a}$ is characterized by a 6D vector describing their relative position at the current timestep: relative distance $d^{ij}$, direction vector $r^{ij}$ under the reference frame of agent $i$, and relative heading vector $h^{ij}=\{\theta^{ij},cos\theta^{ij},sin\theta^{ij}\}$.
@@
$\mathcal{L}_{rep}=\frac{1}{N_{a}}\sum_{i,j,k}\mathrm{max}\{(1-\frac{1}{r}d^{ij}_{k}),0\},$ (15)
where $r$ is the repeller threshold and $d^{ij}_{k}\in\mathbb{R}^{N_{a}\times N_{a}\times(T_{hist}+T_{fut})}$ is the distance between agent $i$ and $j$ at timestep $t$. Practically, the repeller loss serves as valuable prior knowledge.
@@ V Experiments
In this section, we exhibit the capacity and flexibility of our framework by adapting our model to four distinct tasks without retraining: basic trajectory prediction, controllable generation, prediction with noisy history, and prediction with incomplete history. We present two versions of our proposed framework: the foundational GFTD and its variant, GFTD-RePaint, which is specialized for the basic trajectory prediction task.
@@ V-B Trajectory Prediction and Controllable Generation
TABLE IV: Prediction with Incomplete History Data
@@
25% Joint AgentFormer [5] 0.558 / 0.901 0.214 / 0.377 0.624 / 1.286 0.336 / 0.653 0.303 / 0.614 0.407 / 0.766
Ours 0.524 / 0.910 0.174 / 0.294 0.662 / 1.321 0.341 / 0.670 0.314 / 0.627 0.403 / 0.764
50% Joint AgentFormer [5] 0.619 / 1.022 0.219 / 0.378 0.662 / 1.343 0.374 / 0.721 0.327 / 0.649 0.440 / 0.823
Ours 0.511 / 0.897 0.176 / 0.302 0.676 / 1.348 0.347 / 0.677 0.321 / 0.642 0.406 / 0.773
75% Joint AgentFormer [5] 0.647 / 1.034 0.274 / 0.444 0.752 / 1.488 0.440 / 0.830 0.381 / 0.733 0.499 / 0.906
Ours 0.519 / 0.885 0.193 / 0.333 0.735 / 1.441 0.357 / 0.696 0.346 / 0.683 0.430 / 0.808
@@ V-C Prediction with Noisy History
To assess the robustness of our proposed method against noisy input, we perturbed the observed history trajectories by introducing random Gaussian noise $n\sim\mathcal{N}(0,\sigma^{2}I)$ of varying standard deviation $\sigma$. Specifically, we introduced two levels of perturbation: a slight level, characterized by a standard Gaussian noise with a 0.05-meter $\sigma$, and a heavy level, with a standard Gaussian noise with a 0.15-meter $\sigma$. Subsequently, we compared the performance of our model with the Joint AgentFormer baseline under identical experimental conditions.
As shown in TABLE LABEL:table:_noisy_data, we observed that our model performs better at heavy noise levels, providing trustworthy prediction even if the data is highly corrupted. The average JADE/JFDE performance of our method on 5 benchmarks is close to the baseline but only degenerated by $46.58\%/38.38\%$ at the heavy level. By contrast, the performance of the baseline method had degenerated by $57.08\%/50.36\%$ at heavy noise level. To reasonably state that the guided posterior sampling accounted for the robustness against noisy data, we further compared it with GFTD-RePaint. GFTD-RePaint includes gradually replacing the history part of the intermediate noisy sample by ground truth history observation, thus ensuring the generated samples were strictly conditioned on the noisy history data. The experimental results in Fig. 4 show that despite having a near performance to GFTD model under clean data conditions, the GFTD-RePaint performs much worse on noisy data, especially under heavy-level noise. It shows the superiority of our proposed framework in noisy conditions rooted in the design of ”soft conditioning”.
@@ V-D Prediction with Incomplete History
Due to obstacles or sensor failures, historical data may occasionally contain unexpected missing frames. To evaluate model robustness against incomplete data, we randomly select and mask $25\%$, $50\%$, and $75\%$ of the historical trajectory frames, while ensuring that the current frame is always retained. With a known random mask, we only consider the available historical frames when calculating the guidance loss. We employ the same Joint AgentFormer model as our baseline. To ensure a fair comparison, we generate an attention mask for each AgentFormer sub-module based on the known data mask and replace the masked frames with zero values. For both models, we directly re-used the same pre-trained versions as in the other two experiments. As shown in TABLE LABEL:table:_missing_data, GFTD demonstrates competitive performance under conditions of incomplete data. Even when $75\%$ of the historical observations are missing, the average JADE/JFDE performance only declines by $8.9\%$ and $7.4\%$, respectively, still surpassing S-GAN with complete data input by a significant margin. In contrast, Joint AgentFormer fails to produce plausible predictions under high missing data rates.
@@ VI Conclusion & Discussion
In this work, we present the Guided Full Trajectory Diffuser, a novel framework for representing the joint distribution of trajectories leveraging diffusion models, converting trajectory prediction and controllable generation into a unified inverse problem. We formulate the prediction task as spatial-temporal inpainting, a general noisy inverse problem that can be solved through diffusion posterior sampling. Under this framework, the generated trajectories are not rigidly constrained by their historical observations; instead, we gradually enforce such conditions by adjustable posterior guidance. Such a unique design enables flexibility, resulting in not only competitive performance in joint trajectory prediction tasks but also generalizable to scenarios with noise perturbation or incomplete historical data. Moreover, our framework is compatible with plug-and-play modules and various guidance methods, thus expandable to task-oriented enhancements. However, we point out that our current implementation only considers raw trajectories as guidance reference, how to more efficiently exploit the interaction to further improve guidance quality is a topic we will keep working on.